# Ergodic behaviors in reversible 3-state cellular automata

### v1

@@ 4.4 Correlation functions
To characterize the transport phenomena, we compute dynamical two-point of various observables. In particular, we compute correlators of all single-site observables. This space is constrained by the fact that $[\varnothing]_{x}+[+]_{x}+[-]_{x}\equiv\mathds{1}_{x}$, which means that one has to check correlators of only two independent observables. Hence, we consider connected correlators of the charge $c_{x}:=[+]_{x}-[-]_{x}$ and density $d_{x}\equiv c_{x}^{2}=[+]_{x}+[-]_{x}$ operators sampled according to a maximum entropy (infinite temperature) distribution,
@@
[…]
@@
For example, dynamics exhibiting diffusion with the diffusion constant $D$ has either of the two dynamical exponents equal to $21/z=1/$ 2, and the profile of the correlator is Gaussian:
@@
The second common case is ballistic transport $z=1$, which is typically associated with charge-carrying ‘free’ quasiparticles. Another option is the so-called superdiffusive transport with $11/2<1/z<$ 1. A notable example is the Kardar-Parisi-Zhang-type (or KPZ-type) scaling with $31/z=2/$ 3 which, in the Hamiltonian (non-noisy/non-dissipative) one-dimensional lattice systems, occurs in two different known contexts: (i) in anharmonic oscillator chains like the Fermi-Pasta-Ulam systems with exactly three coupled conserved quantities, where the mechanism is explained within the theory of non-linear fluctuating hydrodynamics[32, 33], and (ii) in integrable systems with non-abelian global symmetries [34, 35, 36, 37, 38, 39], where the precise mechanism of the KPZ scaling is still unclear. Finally, the so-called subdiffusive scaling has $20<1/z<1/$ 2 (see, e.g., Refs. [40, 41, 42, 43]). The distribution of all empirically determined dynamical exponents is shown in Figure 8.
Figure 8: Histograms of all dynamical exponents $1/z_{C}$ and $1/z_{D}$ for Classes II, III and IV. The vertical dashed lines, from left to right, represent the values of $1/z\in\{\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$. The counting only includes the physically distinct RCA (with ‘duplicates’ removed).
@@ 5 Class I
$Z^{(r)}_{\ell}=\left<u^{(r,\ell)},u^{(r,\ell)}\right>,\quad\ell=1,2\ldots r\,,$ (45)
@@ 6 Class II
Figure 12: Histograms of all dynamical exponents $1/z_{C}$ and $1/z_{D}$ for sublcasses of Class II. The vertical dashed lines, from left to right, represent the values of $1/z\in\{\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$. Note again that in all such histograms, the counting only includes the physically distinct RCA models with their ‘duplicates’ removed.
@@ 6.1 Subclass IIa
This subclass contains models with a constant number of charges as the range increases. Based on our definition of integrability, such models can be regarded as chaotic. In contrast to the chaotic rules from Class I, which have no conserved charges, models in Subclass IIa have at least one local or quasilocal charge. The presence of charges constrains the decay of correlation functions of local observables that overlap with the charge density. For the rules in Class IIa, where the correlators of range-1 observables decay exponentially, we find range-2 observables that exhibit algebraic decay of the corresponding correlators due to the overlap with charge density. Here, we discuss a number of interesting physical properties of these models, in particular, the existence of quasilocal charges, superdiffusive and subdiffusive transport, and exponential decay of range-1 observables.
@@
The spectrum of the transfer matrix for this model was plotted above in Figure 7. It shows that the rule has three local conserved quantities and no quasilocal charges. The rule has an extensive charge with a density of support 1: $Q_{1}=\sum_{x}[+]_{x}-[-]_{x}$. Indeed, as one can simply infer from the elementary update rule, the difference between the number of red and blue cells is conserved. The correlation function $\hat{C}(x,t)$ then reveals that the transport of the charge $Q_{1}$ is superdiffusive with the KPZ-like dynamical exponent $31/z_{C}\approx 2/$ 3 (see Figure 14).
Figure 14: $\mathcal{CP}+\mathcal{PT}$-symmetric cellular automaton with rule 35216478. (a) $\hat{C}(x,t)$. (b) Cross-sections of $\hat{C}(x,t)$ at increasing fixed times $t$ denoted in the legend. The red dashed line shows the best fit of the Prahofer-Spohn scaling funciton $f_{PS}$, while the black dashed line shows a fit of $\mathrm{exp}(-c|x|^{3})$ with $c\approx 0.2621\pm 10^{-5}.$
In Figure 14, We also compare $\hat{C}(x,t)$ at increasing times $t$ to the Prahofer-Spohn scaling function $f_{PS}$ [52], which is associated with the scaling of correlation function for models in the KPZ universality class. We find that this function does not fit well to our computed correlator. We observe a scaling function that is significantly flatter at $x=0$ and decays more rapidly with increasing $x$ when compared to either Gaussian or KPZ profiles.
@@
Subdiffusive Transport. Some chaotic models in Class IIa exhibit a slower decay of correlators than the diffusive $21/z=1/$ 2. An example of such a subdiffusive model is the cellular automaton with rule 45621378. This model is $\mathcal{P}$ and $\mathcal{C}\mathcal{T}$ symmetric.
The model has three local extensive charges: two with a density of support 1 and one with a support of 4. From the local update rule, two independent charge densities of range 1 can be identified, namely, $q_{1}^{(1)}=[+]-[-]$ and $q_{2}^{(1)}=[\varnothing]$. These conserved quantities correspond to the conservation of the total number of each ‘particle’ species. [^4] While we have three such extensive charges with densities [+],[]delimited-[]delimited-[][+],[-][ + ] , [ - ], and []delimited-[][$\varnothing$][ ∅ ], due to the condition $[+]+[-]+[\varnothing]\equiv\mathds{1}$, only two of them are independent To investigate the transport of these charges, we calculate the correlation function of their densities. In Figure 16 we demonstrate the transport of the charges.
Figure 16: Correlation functions of the subdiffusive $\mathcal{CT+P}$-symmetric cellular automaton with rule 45621378. (a) $\hat{C}(x,t)$ of the observable [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ], (b) Cross-sections of $\hat{C}(x,t)$ at increasing $t$. The inset shows the $\hat{C}(0,t)\sim t^{-1/z}$, with the dashed green line showing the best fit $0.331/z_{C}=$ 0.33. (c) $\hat{D}(x,t)$ of the observable []delimited-[][$\varnothing$][ ∅ ]. (d) Cross-sections of $\hat{D}(x,t)$ at increasing $t$, with the inset showing the decay of the maximum over $x$.
@@
Subdiffusion is often associated with the conservation of dipole moments of some conserved charge density or higher multipole moments [53, 54, 55]. However, the conservation of a dipole moment typically results in a larger dynamical exponent $z=4$. Subdiffusive transport has also been found in different spin chain models [56, 57, 58], where the interplay between interactions, disorder, and boundary effects plays a crucial role in the observed dynamics. Understanding the various microscopic mechanisms responsible for the distinct subdiffusive transport remains an open question.
@@
$[0]=[\varnothing]+[+]+[-]\equiv\mathds{1},\quad[1]=[+]-[-],\quad[2]=-2[\varnothing]+[+]+[-].$ (47)
@@
Figure 17: $\mathcal{CPT}$-symmetric cellular automaton with rule 13825476. (a) $\hat{C}_{3}(x,t)=\langle[11]_{x}(t)[11]_{0}(0)\rangle_{\infty}^{C}$ of an observable with support 2. (b) Cross-sections of $\hat{C}_{3}(x,t)$ at increasing times. The inset shows $C_{3}(t)\sim t^{1/z}$ with the dashed green line showing the best fit of $0.781/z=$ 0.78.
@@ 6.2 Subclass IIb
This subclass contains models with a linearly growing number of local charges, $N_{Q}\sim r$ [^5] We remark that, due to a finite accessible support $r$, such behavior can only be conjectured. Due to the presence of an extensive number of local conserved quantities, we will refer to these models as integrable, however, their local rules $U$ do not satisfy the set-theoretic Yang-Baxter equation and a possible theoretical framework for studying their integrability remains to be understood. Regardless of the presence of conserved charges, the average return time for all integrable rules increases exponentially with the system size.
@@
One of the simplest charges is given by translations of the local density [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ]. The transport of this charge is depicted in Figure 18, where a spreading Gaussian peak associated with diffusion is observed.
Figure 18: Integrable $\mathcal{CP+PT}$-symmetric cellular automaton with rule 26534178 with diffusive transport. (a) $\hat{C}(x,t)$ of the density of charge [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ]. (b) Cross-sections of $\hat{C}(x,t)$ at increasing times. Note the rescaling of the axes by $t^{1/2}$. The insets show $\hat{C}(x=0,t)$ with the best fitting dynamical exponent $0.491/z=$ 0.49.
@@
Figure 19: Correlation functions of the integrable $\mathcal{CP+PT}$-symmetric cellular automaton with rule 25314678. (a) $\hat{C}(x,t)$ of the observable [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ], (b) Cross-sections of $\hat{C}(x,t)$ at increasing $t$. The inset shows the $C(t)\sim t^{-1/z}$, with the dashed green line showing the best fit $0.551/z_{C}=$ 0.55. (c) $\hat{D}(x,t)$ of the observable []delimited-[][$\varnothing$][ ∅ ], which is a density of a conserved charge of this model. (d) Cross-sections of $\hat{D}(x,t)$ at increasing $t$, with the inset showing the decay of the maximum over $x$ with dynamical exponent $0.611/z_{D}=$ 0.61.
@@ 6.3 Subclass IIc
[…]
@@ 7.2 Subclass IIIb
Models in Subclass IIIb have an exponential number of charges and can therefore be considered as superintegrable. For certain models (which are $\mathcal{CPT}$ symmetric), the correlator $C(t)$ decays algebraically, while $D(t)$ exhibits a constant plateau in the thermodynamic limit for all rules. The histogram of all dynamical exponents is shown again in Figure 25, with most models exhibiting values close to $21/z_{C}=1/$ 2. In these models, discussing transport and drawing conclusions from correlators with support 1 turns out to be difficult. The reason is that most of the charges in this class are gliders [59], i.e., their densities $q$ are rigidly translating either left or right, $\mathbb{U}q=\mathbb{T}^{\pm 2}q$.
Figure 25: Histogram of all dynamical exponents $1/z_{C}$ found in Class III. Importantly, all of these result correspond to models in Subclass IIIb. The vertical dashed lines, from left to right, represent the values of $1/z_{C}\in\{\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$.
@@
The system exhibits an algebraic decay of the correlator $C(x,t)$ (see Figure 26), While $D(x,t)$ does not decay to zero. Importantly, however, this rule does not conserve $\sum_{x}\leavevmode\nobreak\ [+]_{x}-[-]_{x}$, so the correlation function $C(x,t)$ is not directly related to transport. We propose that the algebraic decay arises from the overlap between range-1 observables and charges of larger support, a point we will examine further below.
@@
One finds that the simplest charge of range 2 is given by the local density $[20]_{x}$. At range 4, the two charges are translations of local densities $[2000]_{x}$ and $[2020]_{x}$. At every higher range, there is a set of charges, which span all possible combinations of $[20]$ and $[00]$ conspiring to give a basis element of support $r$, of which there are $2^{(\frac{r}{2}-1)}$. All charges constructed in this way are gliders. The densities thus remain perfectly correlated at later times, but they have a constant drift. We propose that the diffusive-like decay of the correlation function for the observable [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ] cannot come from overlaps with gliders, but from other charges, which are of range 6 or higher. These charges can be found relatively easily, as long as one recognizes that the space of gliders is completely independent of them. We list the range-6 charges which are not gliders of range 6 here:
@@
The charge $q_{1}^{(6)}$ has non-zero overlap with the observable $[1]\equiv[+]-[-]$, thus, the transport of this charge contributes to the decay of the correlation function $\hat{C}(x,t)$. Studying the transport of this charge would therefore involve eliminating the drift velocity from the correlation functions of range-1 observables, which could exhibit diffusion-like decay. This is beyond the scope of the present work. We postpone such explorations until a later time.
Figure 26: Correlation functions for the $\mathcal{CPT}$-symmetric cellular automaton with rule 34671528. (a) $\hat{C}(x,t)$, (b) Cross-sections of $\hat{C}(x,t)$ for increasing times $t$. The inset shows $C(t)\sim t^{-1/z_{C}}$, where the dashed green line shows the best fit with $0.51/z_{C}=$ 0.5.
@@
Figure 27: Correlation functions of the $\mathcal{CPT}$-symmetric cellular automaton with rule 23658471. (a) $\hat{C}(x,t)$, (b) Cross-sections of $\hat{C}(x,t)$ at increasing times $t$. The inset shows $C(t)\sim t^{1/z}$ with the dashed green line showing the best fit $0.721/z=$ 0.72.
@@
[…]
However, this charge does not overlap with the observable [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ], the correlator of which shows anomalous decay. We examined the charges of this superintegrable model up to range 12 and did not find any non-glider charges other than $q_{3}^{(6)}$ from Eq. (51). Thus, charges up to this range cannot explain the anomalous decay. This suggests that the observable in $\hat{C}(x,t)$ may have an overlap with one of the quasilocal charges of the model. In Figure 28, we show the scaling of the largest 12 real eigenvalues of the transfer matrix as a function of increasing range. We observe that one of the eigenvalues indeed approaches the unit value exponentially. The anomalous correlation of the observable [+][]delimited-[]delimited-[][+]-[-][ + ] - [ - ] can therefore be a signature of the anomalous transport of one of the quasilocal charges with which it overlaps. However, details of this analysis will also be left to future works.

### v2

@@ 4.4 Correlation functions
To characterize the transport phenomena, we compute dynamical two-point of various observables. In particular, we compute correlators of all single-site observables. This space is constrained by the fact that $\varnothing]_{x}+[+]_{x}+[-]_{x}\equiv\mathds{1}_{x}$, which means that one has to check correlators of only two independent observables. Hence, we consider connected correlators of the charge $c_{x}:=[+]_{x}-[-]_{x}$ and density $d_{x}\equiv c_{x}^{2}=[+]_{x}+[-]_{x}$ operators sampled according to a maximum entropy (infinite temperature) distribution,
@@
$C(t)\overset{\text{def}}{=}\max\limits_{x\in[1,L]}\hat{C}(x,t),\quad D(t)\overset{\text{def}}{=}\max\limits_{x\in[1,L]}\hat{D}(x,t).$ (37)
@@
For example, dynamics exhibiting diffusion with the diffusion constant $D$ has either of the two dynamical exponents equal to $1/z=1/2$, and the profile of the correlator is Gaussian:
@@
The second common case is ballistic transport $z=1$, which is typically associated with charge-carrying ‘free’ quasiparticles. Another option is the so-called superdiffusive transport with $1/2<1/z<1$. A notable example is the Kardar-Parisi-Zhang-type (or KPZ-type) scaling with $1/z=2/3$ which, in the Hamiltonian (non-noisy/non-dissipative) one-dimensional lattice systems, occurs in two different known contexts: (i) in anharmonic oscillator chains like the Fermi-Pasta-Ulam systems with exactly three coupled conserved quantities, where the mechanism is explained within the theory of non-linear fluctuating hydrodynamics[32, 33], and (ii) in integrable systems with non-abelian global symmetries [34, 35, 36, 37, 38, 39], where the precise mechanism of the KPZ scaling is still unclear. Finally, the so-called subdiffusive scaling has $0<1/z<1/2$ (see, e.g., Refs. [40, 41, 42, 43]). The distribution of all empirically determined dynamical exponents is shown in Figure 8.
Figure 8: Histograms of all dynamical exponents $1/z_{C}$ and $1/z_{D}$ for Classes II, III and IV. The vertical dashed lines, from left to right, represent the values of $1/z\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$. The counting only includes the physically distinct RCA (with ‘duplicates’ removed).
@@ 5 Class I
$Z^{(r)}_{\ell}=\sum_{\ell=r}\left<u^{(r,\ell)},u^{(r,\ell)}\right>,\quad\ell=1,2\ldots r\,$, (45)
@@ 6 Class II
Figure 12: Histograms of all dynamical exponents $1/z_{C}$ and $1/z_{D}$ for sublcasses of Class II. The vertical dashed lines, from left to right, represent the values of $1/z\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$. Note again that in all such histograms, the counting only includes the physically distinct RCA models with their ‘duplicates’ removed.
@@ 6.1 Subclass IIa
This subclass contains models with a constant number of charges as the range increases. Based on our definition of integrability, such models can be regarded as chaotic. In contrast to the chaotic rules from Class I, which have no conserved charges, models in Subclass IIa have at least one local or quasilocal charge. The presence of charges constrains the decay of correlation functions of local observables that overlap with the charge density. For the rules in Class IIa, where the correlators of range-1 observables decay exponentially, we find range-2 observables that exhibit algebraic decay of the corresponding correlators due to the overlap with charge density. Here, we discuss a number of interesting physical properties of these models, in particular, the existence of quasilocal charges, superdiffusive and subdiffusive transport (for some models with rather unusual dynamical exponent $z=3$), and exponential decay of range-1 observables.
@@
The spectrum of the transfer matrix for this model was plotted above in Figure 7. It shows that the rule has three local conserved quantities and no quasilocal charges. The rule has an extensive charge with a density of support 1: $Q_{1}=\sum_{x}[+]_{x}-[-]_{x}$. Indeed, as one can simply infer from the elementary update rule, the difference between the number of red and blue cells is conserved. The correlation function $\hat{C}(x,t)$ then reveals that the transport of the charge $Q_{1}$ is superdiffusive with the KPZ-like dynamical exponent $1/z_{C}\approx 2/3$ (see Figure 14).
Figure 14: $\mathcal{CP}+\mathcal{PT}$-symmetric cellular automaton with rule 35216478. (a) $\hat{C}(x,t)$. (b) Cross-sections of $\hat{C}(x,t)$ at increasing fixed times $t$ denoted in the legend. The red dashed line shows the best fit of the Prahofer-Spohn scaling function $f_{PS}$, while the black dashed line shows a fit of $\mathrm{exp}(-c|x|^{3})$ with $c\approx 0.2621\pm 10^{-5}.$
In Figure 14, we compare $\hat{C}(x,t)$ at increasing times $t$ to the Prähofer–Spohn scaling function $f_{PS}$ [52], which is associated with the scaling of the correlation function in models belonging to the KPZ universality class. We find that this function does not fit well with our computed correlator. We observe a scaling function that is significantly flatter at $x=0$ and decays more rapidly with increasing $x$ than either the Gaussian or KPZ profiles.
@@
Because our simulations have so far been done using a relatively small number of time steps (up to $t=4096$ for systems of similar size), here, we perform a more extensive simulation to confirm the existence of the anomalous dynamical exponent $z=3/2$. In particular, we compute the peak of the correlation function $\hat{C}(0,t)$ for times up to $t=2^{19}$ and for a system with size $2^{20}$. The results are shown in Figure 16, confirming that the model also exhibits the dynamical exponent $z=3/2$ at late times.
Figure 16: Extended simulation of the CA with rule 35216478. The decay of the correlator at $x=0$ is best fit with $1/z=0.656\pm 0.005$. The dashed gray line indicates a decay of $t^{-2/3}$ to guide the eye.
Subdiffusive Transport with $z=3$. Some chaotic models in Class IIa exhibit a slower decay of correlators than the diffusive $1/z=1/2$. An example of such a subdiffusive model is the cellular automaton with rule 45621378. This model is $\mathcal{P}$ and $\mathcal{C}\mathcal{T}$ symmetric.
The model has three local extensive charges: two with a density of support 1 and one with a support of 4. From the local update rule, two independent charge densities of range 1 can be identified, namely, $q_{1}^{(1)}=[+]-[-]$ and $q_{2}^{(1)}=[\varnothing]$. These conserved quantities correspond to the conservation of the total number of each ‘particle’ species. To investigate the transport of these charges, we calculate the correlation function of their densities. In Figure 17 we demonstrate the transport of the charges.
Figure 17: Rule 45621378. (a) $\hat{C}(x,t)$ (b) Cross-sections of $\hat{C}(x,t)$ at increasing $t$. The inset shows the $\hat{C}(0,t)\sim t^{-1/z}$, with the dashed green line showing the best fit $1/z_{C}=0.33$. (c) $\hat{D}(x,t)$. (d) Cross-sections of $\hat{D}(x,t)$ at increasing $t$, with the inset showing the decay of the maximum over $x$.
@@
As in the previous example, we examine whether this type of transport persists for asymptotically long times. We perform an extended simulation for a system of size $2^{20}$, up to times of $2^{19}$. The results are shown in Figure 18 confirming that the dynamical exponent $z=3$ remains valid on much longer timescales.
Figure 18: Decay of the correlator $\hat{C}(x,t)$ of the CA with rule 45621378 at $x=0$ with an algebraic best fit $1/z=0.3266\pm 0.0005$. The dashed gray line denotes a decay of $t^{-1/3}$ to guide the eye.
Subdiffusion is often associated with the conservation of dipole moments of some conserved charge density or higher multipole moments [53, 54, 55]. However, the conservation of a dipole moment typically results in a larger dynamical exponent $z=4$. Subdiffusive transport has also been found in different spin chain models [56, 57, 58], where the interplay between interactions, disorder, and boundary effects plays a crucial role in the observed dynamics. We also find a number of models in this class for which the dynamical exponents converge to $z=4$ (see Appendices B and C). Understanding the various microscopic mechanisms responsible for the distinct subdiffusive transport remains an open question.
@@
$0]=[\varnothing]+[+]+[-]\equiv\mathds{1},\quad[1]=[+]-[-],\quad[2]=-2[\varnothing]+[+]+[-].$ (47)
@@
Figure 19: $\mathcal{CPT}$-symmetric cellular automaton with rule 13825476. (a) $\hat{C}_{3}(x,t)=\langle[11]_{x}(t)[11]_{0}(0)\rangle_{\infty}^{C}$ of an observable with support 2. (b) Cross-sections of $\hat{C}_{3}(x,t)$ at increasing times. The inset shows $C_{3}(t)\sim t^{1/z}$ with the dashed green line showing the best fit of $1/z=0.78$.
@@ 6.2 Subclass IIb
This subclass contains models with a linearly growing number of local charges, $N_{Q}\sim r$. [^4] We remark that, due to a finite accessible support $r$, such behavior can only be conjectured. Due to the presence of an extensive number of local conserved quantities, we will refer to these models as integrable, however, their local rules $U$ do not satisfy the set-theoretic Yang-Baxter equation and a possible theoretical framework for studying their integrability remains to be understood. Regardless of the presence of conserved charges, the average return time for all integrable rules increases exponentially with the system size.
@@
One of the simplest charges is given by translations of the local density [+][][+]-[-]. The transport of this charge is depicted in Figure 20, where a spreading Gaussian peak associated with diffusion is observed.
Figure 20: Integrable $\mathcal{CP+PT}$-symmetric cellular automaton with rule 26534178 with diffusive transport. (a) $\hat{C}(x,t)$ of the density of charge [+][][+]-[-]. (b) Cross-sections of $\hat{C}(x,t)$ at increasing times. Note the rescaling of the axes by $t^{1/2}$. The insets show $\hat{C}(x=0,t)$ with the best fitting dynamical exponent $1/z=0.49$.
@@
Figure 21: Correlation functions of the integrable $\mathcal{CP+PT}$-symmetric cellular automaton with rule 25314678. (a) $\hat{C}(x,t)$ of the observable [+][][+]-[-], (b) Cross-sections of $\hat{C}(x,t)$ at increasing $t$. The inset shows the $C(t)\sim t^{-1/z}$, with the dashed green line showing the best fit $1/z_{C}=0.55$. (c) $\hat{D}(x,t)$ of the observable $\varnothing$], which is a density of a conserved charge of this model. (d) Cross-sections of $\hat{D}(x,t)$ at increasing $t$, with the inset showing the decay of the maximum over $x$ with dynamical exponent $1/z_{D}=0.61$.
@@ 6.3 Subclass IIc
Figure 23: Correlators of dynamical charges of the $\mathcal{CT}+\mathcal{P}$-symmetric cellular automaton with rule 45378621. (a) $\hat{C}_{q_{3}}(x,t)$, (b) Cross-sections of $\hat{C}_{q_{3}}(x,t)$, (c) $\hat{C}_{q_{4}}(x,t)$, (d) Cross-sections of $\hat{C}_{q_{4}}(x,t)$. All cross-sections and insets are plotted for times $t\penalty 10000\ {\rm mod}\penalty 10000\ 6=0$ as the correlator oscillates with period 3 due to the period of the dynamical charge. The insets on the plots in the right panel show $\hat{C}_{q_{3}}(0,t)$ and $\hat{C}_{q_{4}}(0,t)$ which decay algebraically $\sim t^{-2/3}$. Best fits are shown with green dashed lines.
@@ 7.2 Subclass IIIb
Models in Subclass IIIb have an exponential number of charges and can therefore be considered as superintegrable. For certain models (which are $\mathcal{CPT}$ symmetric), the correlator $C(t)$ decays algebraically, while $D(t)$ exhibits a constant plateau in the thermodynamic limit for all rules. The histogram of all dynamical exponents is shown again in Figure 27, with most models exhibiting values close to $1/z_{C}=1/2$. In these models, discussing transport and drawing conclusions from correlators with support 1 turns out to be difficult. The reason is that most of the charges in this class are gliders [59], i.e., their densities $q$ are rigidly translating either left or right, $\mathbb{U}q=\mathbb{T}^{\pm 2}q$.
Figure 27: Histogram of all dynamical exponents $1/z_{C}$ found in Class III. Importantly, all of these result correspond to models in Subclass IIIb. The vertical dashed lines, from left to right, represent the values of $1/z_{C}\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{3}{5},\frac{2}{3}\}$.
@@
The system exhibits an algebraic decay of the correlator $C(x,t)$ (see Figure 28), While $D(x,t)$ does not decay to zero. Importantly, however, this rule does not conserve $\sum_{x}\penalty 10000\ [+]_{x}-[-]_{x}$, so the correlation function $C(x,t)$ is not directly related to transport. We propose that the algebraic decay arises from the overlap between range-1 observables and charges of larger support, a point we will examine further below.
@@
One finds that the simplest charge of range 2 is given by the local density $20]_{x}$. At range 4, the two charges are translations of local densities $2000]_{x}$ and $2020]_{x}$. At every higher range, there is a set of charges, which span all possible combinations of $[20]$ and $[00]$ conspiring to give a basis element of support $r$, of which there are $2^{(\frac{r}{2}-1)}$. All charges constructed in this way are gliders. The densities thus remain perfectly correlated at later times, but they have a constant drift. We propose that the diffusive-like decay of the correlation function for the observable [+][][+]-[-] cannot come from overlaps with gliders, but from other charges, which are of range 6 or higher. These charges can be found relatively easily, as long as one recognizes that the space of gliders is completely independent of them. We list the range-6 charges which are not gliders of range 6 here:
@@
The charge $q_{1}^{(6)}$ has non-zero overlap with the observable $1]\equiv[+]-[-]$, thus, the transport of this charge contributes to the decay of the correlation function $\hat{C}(x,t)$. Studying the transport of this charge would therefore involve eliminating the drift velocity from the correlation functions of range-1 observables, which could exhibit diffusion-like decay. This is beyond the scope of the present work. We postpone such explorations until a later time.
Figure 28: Correlation functions for the $\mathcal{CPT}$-symmetric cellular automaton with rule 34671528. (a) $\hat{C}(x,t)$, (b) Cross-sections of $\hat{C}(x,t)$ for increasing times $t$. The inset shows $C(t)\sim t^{-1/z_{C}}$, where the dashed green line shows the best fit with $1/z_{C}=0.5$.
@@
Figure 29: Correlation functions of the $\mathcal{CPT}$-symmetric cellular automaton with rule 23658471. (a) $\hat{C}(x,t)$, (b) Cross-sections of $\hat{C}(x,t)$ at increasing times $t$. The inset shows $C(t)\sim t^{1/z}$ with the dashed green line showing the best fit $1/z=0.72$.
@@
$\begin{split}q^{(6)}_{3}=&\,\,4[202]+4[222]+8[0101]+4[0121]+4[2201]+4[2101]+2[2121]+2[2221]\\ &+2[01212]-4[01022]-2[01222]+2[21012]+4[01012]+2[22012]\\ &+[21212]+[22212]-2[21022]-2[22022]-[21222]-[22222].\end{split}$ (51)
However, this charge does not overlap with the observable [+][][+]-[-], the correlator of which shows anomalous decay. We examined the charges of this superintegrable model up to range 12 and did not find any non-glider charges other than $q_{3}^{(6)}$ from Eq. (51). Thus, charges up to this range cannot explain the anomalous decay. This suggests that the observable in $\hat{C}(x,t)$ may have an overlap with one of the quasilocal charges of the model. In Figure 30, we show the scaling of the largest 12 real eigenvalues of the transfer matrix as a function of increasing range. We observe that one of the eigenvalues indeed approaches the unit value exponentially. The anomalous correlation of the observable [+][][+]-[-] can therefore be a signature of the anomalous transport of one of the quasilocal charges with which it overlaps. However, details of this analysis will also be left to future works.